. To what volume should 100 mL of any weak 0.30 molar monobasic acid (HA) or mono-acidic base (BOH) be diluted in order to triple its percentage ionization?


1
Expert's answer
2021-06-21T05:32:11-0400

Given that a weak monobasic acid (say HA) has a concentration of 0.3M.

We know pH=−log [H+

]=5

Hence, [H+)=10−5

The ionisation reaction is as follows:

HA       ⇋     H+

+A−

(1−α)c       αc     αc

Here, α is the degree of ionisation and c is the concentration of an acid.

The relationship between degree of ionization and concentration is α=c

[H]+

On substituting values in the above equation, we get

α=0.005

10−5​=2×10−3

Hence, the degree of ionisation is 2×10-3

2 × 0.3 × 100

= 60ml